[Hallc_running] [New Logentry] RC update

mack at jlab.org mack at jlab.org
Mon Mar 25 10:15:02 EDT 2019

Logentry Text:
Current Status
For the production setting at Q2 = 2.115, W = 2.95 , data taking at the SHMS/K+ angle of 8.48 deg is as done as it's going to get. Beam went off at 08:30 for beam studies.  

Plans - After the downstream beampipe is radiologically surveyed, the SHMS angle needs to move down to 6.2 degrees (or a bit smaller). Then we continue running on swing shift. Shift leaders keep track of the requested mix of LH2 and 10cm Dummy runs, and pass your completion status down to the next shift. 

Opportunistic accesses during beam studies - Debitwa may try to recover a dead wire in the SHMS . Markus D may give a tour. I'd like to keep the Hall locked since there should be plenty of keys, but whatever the SHMS angle setting crew wants is fine.  

Announcement - We will have an RC meeting today at 4:30 PM in the counting house. Bluejeans Conference ID : 861439920

Kaon LT run plans : https://hallcweb.jlab.org/wiki/index.php/Main_Page#Kaon_LT_experiment
Information for Shift Takers: https://hallcweb.jlab.org/wiki/index.php/Information_for_Shift_Takers


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3671968
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190325/1581f8fb/attachment.html>

More information about the Hallc_running mailing list